HUBBLE SPACE TELESCOPE DAILY REPORT #1376 PERIOD COVERED: 0000Z (UTC) 04/24/95 - 0000Z (UTC) 04/25/95 Daily Status Report as of 115/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of FGS 6147 (Extragalactic Astronomy and Astrophysics Final: Cycle 4 Observations) The FGSs were used to make observations of targets 0637-752IN, 1120-REF, 1121-REF, INCA221-39, NEWEGOD-57, 520-REF, 521-REF, 522-REF and NEWHIP-57. The objective of this long-term program is to determine the rotation of the HIPPARCOS reference relative to the extragalactic frame. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5092 (GTO Low Latitude Parallel Program: Cycle 4) The WF/PC-2 was used to make observations taken in parallel with the prime science observations. This is part of a survey to take pictures of a large number of fields within 30 degrees of the galactic plane. The observations were completed as planned, and no problems were reported. 1.3 Completed Three Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of TOL1220+05 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 5397 (Determination of the Extragalactic Distance Scale: Cycle 4 High) The WF/PC-2 was used to make images of galaxy NGC-4414 to be used in measuring the periods of Cepheid variable stars within this galaxy. The observations were completed as planned, and no anomalies were noted. 1.6 Completed WF/PC-2 6142 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Smaller Set Part II) The WF/PC-2 was used to make Earth Flat calibration images for the ultraviolet and narrow-band filters to correct for the telescope's illumination pattern. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed FOS/BL 6146 (FOS Location of Spectra and Aperture - Part 2) The Faint Object Spectrograph (Blue Side) was used to conduct a test to determine the repeatability of positioning the filter-grating wheel. All observations were internal. A wave length calibration image was also taken. All observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Three Sets of WF/PC-2 5091 (GTO High Latitude Parallel Program: Cycle 4) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations greater than +/-30 degrees. The observations were executed as scheduled, and no problems were reported. 1.9 Completed WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1C: 4-Chip STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 5981 (Calibration of Nearby Type IA supernova as Standard Candles: NGC-4639) The WF/PC-2 was used to make observations of Cepheid variable stars in Galaxy NGC-4639 (which has produced a Type IA Supernova) to calibrate the nova's absolute brightness so that it can be used as a standard candle for distance measurements. The observations were executed as scheduled, and on anomalies were noted. 1.11 Completed GHRS/2 5728 (Absorption Cloud Physics in Seyfert Galaxy Nuclei: Cycle 2) The High Resolution Spectrograph (Side Two) was used to make spectrographic observations of a C IV region in Seyfert galaxy NGC-3516 to study the structure and velocity in the line of sight. Variability of the spectra on a time scale of months has been reported in this object. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Notes: Occasional bit slips while dumping ESTR #2 are still occurring. TALOS, which is monitoring the CoP activity, was down for four hours yesterday. The problem was corrected, and the monitoring is continuing uneventfully. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-